OpenStudy (anonymous):

What is the value of y in the solution to the following system of equations? 2x + y = -4 5x + 3y = -6 8 6 -6 -8

OpenStudy (jdoe0001):

so solve the system of equations :)

OpenStudy (anonymous):

I would if I knew what to do :(

OpenStudy (jdoe0001):

cancel out "x" from both

OpenStudy (anonymous):

What you're going to want to do is eliminate the x's in both equations and keep only the y's. You could either multiply the first equation by 5 on both sides and multiply the second by 2 on both sides and subtract both equations, or isolate x and substitute the value you get for x in either equation.

OpenStudy (jdoe0001):

2x + y = -4 # multiply it by 5 5x + 3y = -6 # multiply it by -2 what would you get?

OpenStudy (anonymous):

still kinda confused

OpenStudy (anonymous):

There you go, and then just subtract the equations. (for instance you would cancel the x's because they would both be 10 - 10 = 0

OpenStudy (anonymous):

ohhh

OpenStudy (anonymous):

Remember to multiply everything by 5! Then, solve either equation for y and there you have it!

OpenStudy (anonymous):

alright thanks

OpenStudy (anonymous):

so its y = 8